# Nonisotropic Gaussian Diffusion for Realistic 3D Human Motion Prediction

### v1

@@ 6 Conclusion
We present SkeletonDiffusion, a new latent diffusion model with an explicit inductive bias on the human skeleton trained with a novel nonisotropic Gaussian diffusion formulation. We achieve state-of-the-art performance on stochastic HMP by generating motions that are simultaneously realistic and diverse while being robust to limb stretching according to the evaluation metrics.
@@
Similar to previous methods, we restrict our experiments to standard human skeletons, where fine-grained joints (e.g., fingers, facial expression) are not considered. Unfortunately, at the present date, such data are scarce and difficult to capture. Applying our non-isotropic diffusion on skeletons with more expressive structures is an inspiring direction for future works. We also believe it may be worth experimenting with different definitions of $\boldsymbol{\Sigma}_{N}$, depending on the applicative context, and eventually exploring non-Gaussian diffusion formulations.
@@ Acknowledgments
This work was supported by the ERC Advanced Grant SIMULACRON. Thanks to Dr. Almut Sophia Koepke, Yuesong Shen and Shenhan Qian for the proofreading and feedback, Lu Sang for the discussion, Stefania Zunino and the whole CVG team for the support.
@@ Appendix B Network architecture
SkeletonDiffusion’s architecture builds on top of Typed-Graph (TG) convolutions [65], a type of graph convolutions designed particularly for human motion prediction. The conditional autoencoder consists of two shallow TG GRU [65]. To obtain a strong temporal representation of arbitrary length, thus fitting both observation and ground truth future, we pass the encoder’s last GRU state to a TG convolutional layer [65]. The denoiser network consists of a custom architecture of stacked residual blocks of TG convolutions and TG Attention layers. Details will be made available through the code release.
@@ Appendix C Training Details
The conditional autoencoder is trained for 300 epochs on AMASS, 200 on FreeMan, and 100 on H36M. To avoid collapse towards the motion mean of the training data [9, 78], we employ curricular learning [8, 78, 1] and learn to reconstruct sequences with random length $l$, sampled from a discrete uniform distribution $l\sim{\mathcal{U}}\{1,\tilde{F}\}$. Specifically, we increase the upper bound of the motion length $\tilde{F}$ to the original future timewindow $F$ after the first 10 epochs with a cosine scheduler. The denoiser network is trained with $T=10$ diffusion steps and a learning rate of 0.005 for 150 epochs. We employ a cosine scheduler [58] for ${\alpha}_{t}$ and implement an exponential moving average of the trained diffusion model with a decay of 0.98. Inference sampling is drawn from a DDPM sampler [31]. Both networks are trained with Adam on PyTorch. The biggest version of our model (AMASS) consists of 34M parameters and is trained on a single NVIDIA GPU A40 for 6 days. For AMASS, we measure an inference time of 471 milliseconds for a single batch on a NVIDIA GPU A40, in line with the latest DM works.
@@ D.1 Metrics in Stochastic HMP
$\mathrm{MMFDE}(\tilde{{\textbf{Y}}},{}^{\mathrm{MM}}{\textbf{Y}})=$ $\min_{(i,j)\in\mathcal{M}}\|\tilde{{\textbf{Y}}}^{i}_{F}-{}^{\mathrm{MM}}{\textbf{Y}}^{j}_{F}\|_{2}$ (56)
@@
Precision Multimodal GT Diversity Realism Body Realism
Type Method mean $\downarrow$ RMSE $\downarrow$
ADE $\downarrow$ FDE $\downarrow$ MMADE $\downarrow$ MMFDE $\downarrow$ APD $\uparrow$ CMD $\downarrow$ FID str jit str jit
GAN HP-GAN [6] 0.858 0.867 0.847 0.858 7.214 - - - - - -
DeLiGAN [30] 0.483 0.534 0.520 0.545 6.509 - - - - - -
VAE TPK [77] 0.461 0.560 0.522 0.569 6.723 6.326 0.538 6.69 0.24 8.37 0.31
DLow [88] 0.425 0.518 0.495 0.531 11.741 4.927 1.255 7.67 0.28 9.71 0.36
GSPS [54] 0.389 0.496 0.476 0.525 14.757 10.758 2.103 4.83 0.19 6.17 0.24
Motron [65] 0.375 0.488 0.509 0.539 7.168 40.796 13.743 - - - -
DivSamp [19] 0.370 0.485 0.475 0.516 15.310 11.692 2.083 6.16 0.23 7.85 0.29
Other STARS [82] 0.358 0.445 0.442 0.471 15.884 - - - - - -
SLD [81] 0.348 0.436 0.435 0.463 8.741 - - - - - -
DM MotionDiff [80] 0.411 0.509 0.508 0.536 7.254 - - 8.04 0.59 10.21 0.77
HumanMAC [16] 0.369 0.480 0.509 0.545 6.301 - - 4.01 0.46 6.04 0.57
BeLFusion [5] 0.372 0.474 0.473 0.507 7.602 5.988 0.209 5.39 0.17 6.63 0.22
CoMusion [67] 0.350 0.458 0.494 0.506 7.632 3.202 0.102 4.61 0.41 5.97 0.56
DM SkeletonDiff 0.344 0.450 0.487 0.512 7.249 4.178 0.123 3.90 0.16 4.96 0.21
Table 5: Comparison on Human3.6M. Bold and underlined results correspond to the best and second-best results, respectively.
@@ D.3 Datasets
As mentioned in the main paper, often metrics hide or may be influenced by artifacts. Inspecting qualitative results can lead to better insights into the effective SHMP methods’ performance. Previous works [5, 67, 16, 77, 88, 54, 19] visualize the diversity of the predictions by overlapping the skeleton of multiple motions in different colors. This representation is limited and not well suited to identify motion irregularities. We propose to fit a SMPL mesh to each skeleton pose to ease inspection of the results, while preserving the semanticity of the prediction. Ill-posed predictions can thus be easily spotted through the erroneous SMPL fitting. For completeness, we still report the historical visualizations in Appendix G.
@@
Norm Type mean $\downarrow$ RMSE $\downarrow$
Frob 0.480 0.539 0.561 0.575 9.468 12.066 3.26 0.20 4.54 0.26
Appendix E Additional Experiments
E.1 Human3.6M
E.2 Diversity and Body Realism
Precision Multimodal GT Diversity Realism Body Realism
Base of $\boldsymbol{\Sigma}_{N}$ mean $\downarrow$ RMSE $\downarrow$
$\mathbf{R}$ 0.481 0.540 0.562 0.574 9.504 11.542 3.16 0.20 4.51 0.27
$\mathbf{R}^{hip}$ 0.475 0.543 0.558 0.579 8.629 12.499 3.14 0.19 4.35 0.25
Appendix F Further Analysis
F.1 On Nonisotropic and Isotropic Diffusion Training
F.2 Discussion on Correlation Matrix $\boldsymbol{\Sigma}_{N}$
@@
F.3 Correlations of Latent Space
We visualize the skeleton latent space in terms of the correlation among different latent joint dimensions. To this end, we embed all AMASS test segments in the latent space, and compute the first principal component along the each joint dimension separately. For each embedding, we then plot the principal component of two joint dimensions against each other. In Fig. 12, we show 50 random test segments and for each 15 diffused latents. Our latent space reflects correlations connected body joints that are expected (e.g. LHip and RHip) or are less intuitive (e.g. Neck and Hip always show in the same space direction), while other joints do not exhibit univocal correlations (e.g. Wrist and Ankle of the same body side). Weak correlations (probably related to the walking pattern) can be observed between opposite joints of the lower and upper body such as RHip and LElbow.

### v2

@@ 6 Conclusion
We present SkeletonDiffusion, a latent diffusion model with an explicit inductive bias on the human skeleton trained with a novel nonisotropic Gaussian diffusion formulation. We achieve state-of-the-art performance on stochastic HMP by generating motions that are simultaneously realistic and diverse while being robust to limb stretching according to the evaluation metrics.
@@
Similar to previous methods, we restrict our experiments to standard human skeletons, without considering fine-grained joints (e.g., fingers, facial expression). Unfortunately, such data are scarce and difficult to capture. While our body realism metrics address previously disregarded aspects, evaluating stochastic HMP and particularly diversity remains an open challenge.
@@ Acknowledgments
This work was supported by the ERC Advanced Grant “SIMULACRON” (agreement #884679), GNI Project “AI4Twinning”, and DFG project CR 250/26-1 “4D YouTube”. Thanks to Dr. Almut Sophia Koepke, Yuesong Shen and Shenhan Qian for the proofreading and feedback, Lu Sang for the discussion, Jialin Yang for the applications, Stefania Zunino and the whole CVG team for the support.
@@ Appendix B Network architecture
SkeletonDiffusion’s architecture builds on top of Typed-Graph (TG) convolutions [67], a type of graph convolutions designed particularly for human motion prediction. The conditional autoencoder consists of two shallow TG GRU [67]. To obtain a strong temporal representation of arbitrary length, thus fitting both observation and ground truth future, we pass the encoder’s last GRU state to a TG convolutional layer [67]. The denoiser network consists of a custom architecture of stacked residual blocks of TG convolutions and TG Attention layers. Details are available through the code implementation.
@@ Appendix C Training Details
The conditional autoencoder is trained for 300 epochs on AMASS, 200 on FreeMan, and 100 on H36M. In the autoencoder training, to avoid collapse towards the motion mean of the training data [9, 80], we employ curricular learning [8, 80, 1] and learn to reconstruct sequences with random length $l$, sampled from a discrete uniform distribution $l\sim{\mathcal{U}}\{1,\tilde{F}\}$. Specifically, we increase the upper bound of the motion length $\tilde{F}$ to the original future timewindow $F$ after the first 10 epochs with a cosine scheduler. The denoiser network is trained with $T=10$ diffusion steps and a learning rate of 0.005 for 150 epochs. We employ a cosine scheduler [59] for ${\alpha}_{t}$ and implement an exponential moving average of the trained diffusion model with a decay of 0.98. Inference sampling is drawn from a DDPM sampler [31]. Both networks are trained with Adam on PyTorch. The biggest version of our model (AMASS) consists of 34M parameters and is trained on a single NVIDIA GPU A40 for 6 days. For AMASS, we measure an inference time of 471 milliseconds for a single batch on a NVIDIA GPU A40, in line with the latest DM works.
Precision Multimodal GT Diversity Realism Body Realism
Base of $\boldsymbol{\Sigma}_{N}$ mean $\downarrow$ RMSE $\downarrow$
$\mathbf{R}$ 0.481 0.540 0.562 0.574 9.504 11.542 3.16 0.20 4.51 0.27
$\mathbf{R}^{hip}$ 0.475 0.543 0.558 0.579 8.629 12.499 3.14 0.19 4.35 0.25
@@ D.1 Metrics in Stochastic HMP
$\mathrm{MMFDE}(\tilde{\textbf{Y}},{}^{\mathrm{MM}}\textbf{Y})=$ $\min_{(i,j)\in\mathcal{M}}\|\tilde{\textbf{Y}_{F}}^{i}-{}^{\mathrm{MM}}\textbf{Y}_{F}^{j}\|_{2}$ (56)
@@
The Fréchet inception distance (FID) is computed for H36M only (as in [4, 13, 62]), as obtaining the necessary classifier to compute the features is not trivial: AMASS does not have class labels (recently, BABEL [62] annotated only 1% of the test data), and FreeMan’s annotations do not map into specific classes.
@@ D.2 Baselines
Precision Multimodal GT Diversity Realism Body Realism
Norm Type mean $\downarrow$ RMSE $\downarrow$
Frob 0.480 0.539 0.561 0.575 9.468 12.066 3.26 0.20 4.54 0.26
@@
Precision Multimodal GT Diversity Realism Body Realism
mean $\downarrow$ RMSE $\downarrow$
w/o-TG-Att 0.502 0.567 0.576 0.597 8.021 14.934 3.90 0.20 5.31 0.27
iso 0.499 0.553 0.568 0.583 8.788 15.603 3.72 0.18 4.93 0.24
noniso 0.489 0.547 0.567 0.581 9.483 11.812 2.77 0.20 4.06 0.27
Ours (SkeletonDiffusion) 0.480 0.545 0.562 0.579 9.456 11.418 3.15 0.20 4.45 0.26
Table 7: Ablations on the AMASS dataset [53].
Precision Multimodal GT Diversity Realism Body Realism
mean $\downarrow$ RMSE $\downarrow$
Ours+Past 0.574 0.584 0.607 0.599 9.856 16.993 10.16 0.24 11.04 0.38
Ours+DCT 0.534 0.572 0.595 0.600 11.215 16.783 5.20 0.25 7.59 0.35
Ours (SkeletonDiffusion) 0.480 0.545 0.562 0.579 9.456 11.418 3.15 0.20 4.45 0.26
Table 8: Additional ablations on AMASS [53] for discarded components.
@@ D.3 Datasets
As mentioned in the main paper, often metrics hide or may be influenced by artifacts. Inspecting qualitative results can lead to better insights into the effective SHMP methods’ performance. Previous works [5, 69, 16, 79, 91, 55, 19] visualize the diversity of the predictions by overlapping the skeleton of multiple motions in different colors. This representation is limited and not well suited to identify motion irregularities. We propose to fit a SMPL mesh to each skeleton pose to ease inspection of the results, while preserving the semanticity of the prediction. Ill-posed predictions can thus be easily spotted through the erroneous SMPL fitting. For completeness, we still report the historical visualizations in Fig. 8.
Appendix E Further Analysis
E.1 Correlations of Latent Space
We visualize the latent space in terms of the correlation among different latent joint dimensions. To this end, we embed all AMASS test segments in the latent space, and compute the first principal component along the each joint dimension separately. For each embedding, we then plot the principal component of two joint dimensions against each other. In Fig. 9, we show 50 random test segments and for each 15 diffused latents. Our latent space reflects correlations connected body joints that are expected (e.g. LHip and RHip) or are less intuitive (e.g. Neck and Hip always show in the same space direction), while other joints do not exhibit univocal correlations (e.g. Wrist and Ankle of the same body side). Weak correlations (probably related to the walking pattern) can be observed between opposite joints of the lower and upper body such as RHip and LElbow.
@@ E.1 Correlations of Latent Space
Type Method mean $\downarrow$ RMSE $\downarrow$
ADE $\downarrow$ FDE $\downarrow$ MAE $\downarrow$ MMADE $\downarrow$ MMFDE $\downarrow$ APD $\uparrow$ CMD $\downarrow$ FID str jit str jit
Alg Zero-Velocity 0.597 0.884 6.753 0.683 0.909 0.000 22.812 0.606 0.00 0.00 0.00 0.00
GAN HP-GAN [6] 0.858 0.867 - 0.847 0.858 7.214 - - - - - -
DeLiGAN [30] 0.483 0.534 - 0.520 0.545 6.509 - - - - - -
VAE TPK [79] 0.461 0.560 8.056 0.522 0.569 6.723 6.326 0.538 6.69 0.24 8.37 0.31
DLow [91] 0.425 0.518 6.856 0.495 0.531 11.741 4.927 1.255 7.67 0.28 9.71 0.36
GSPS [55] 0.389 0.496 7.171 0.476 0.525 14.757 10.758 2.103 4.83 0.19 6.17 0.24
Motron [67] 0.375 0.488 - 0.509 0.539 7.168 40.796 13.743 - - - -
DivSamp [19] 0.370 0.485 6.257 0.475 0.516 15.310 11.692 2.083 6.16 0.23 7.85 0.29
Other STARS [84] 0.358 0.445 - 0.442 0.471 15.884 - - - - - -
SLD [83] 0.348 0.436 - 0.435 0.463 8.741 - - - - - -
DM MotionDiff [82] 0.411 0.509 - 0.508 0.536 7.254 - - 8.04 0.59 10.21 0.77
HumanMAC [16] 0.369 0.480 6.167 0.509 0.545 6.301 - - 4.01 0.46 6.04 0.57
BeLFusion [5] 0.372 0.474 6.107 0.473 0.507 7.602 5.988 0.209 5.39 0.17 6.63 0.22
CoMusion [69] 0.350 0.458 5.904 0.494 0.506 7.632 3.202 0.102 4.61 0.41 5.97 0.56
DM SkeletonDiff 0.344 0.450 5.556 0.487 0.512 7.249 4.178 0.123 3.90 0.16 4.96 0.21
Table 9: Comparison on Human3.6M [34]. Bold and underlined results correspond to the best and second-best results, respectively.
E.2 Discussion on Correlation Matrix $\boldsymbol{\Sigma}_{N}$
@@
E.3 On the Convergence of Nonisotropic Diffusion
E.4 Ablations of SkeletonDiffusion
In Tab. 7, we report the ablations discussed in Sec. 5.2. We compare the effect of TG-Attention layers on isotropic diffusion ($\boldsymbol{\Sigma}_{N}=I$ and ${\gamma}_{t}=0$) and analyze nonisotropic diffusion with our covariance reflecting joint connections $\boldsymbol{\Sigma}_{N}$ (Eq. 4) in the variant where ${\gamma}_{t}=1$ (as in Eq. 3) and our blending with the scheduler ${\gamma}_{t}$ (Eq. 5). Further experiments, such as fine-tuning the encoder responsible for embedding the past observation or representing motion data via the Discrete Cosine Transform (DCT) [16] are reported in Tab. 8. From the low precision results of the latter experiment and referring to Tab. 1, we speculate that while DCT seems suitable for transformer-based diffusion models operating in input space [69, 16], extracting features directly from Euclidean motion space seems a better choice for latent diffusion models (BeLFusion [5] and our method).
Appendix F Additional Experiments
Precision Multimodal GT Diversity Realism Body Realism
Type Method mean $\downarrow$ RMSE $\downarrow$
ADE $\downarrow$ FDE $\downarrow$ MAE $\downarrow$ MMADE $\downarrow$ MMFDE $\downarrow$ APD $\uparrow$ CMD $\downarrow$ str jit str jit
Alg ZeroVelocity 0.764 1.016 10.921 0.785 1.019 0.000 40.695 4.52 0.00 4.52 0.00
VAE DLow 0.596 0.652 9.188 0.615 0.654 13.776 12.754 8.79 0.43 11.73 0.63
DivSamp 0.583 0.690 10.758 0.617 0.698 23.878 46.594 12.38 0.82 18.11 1.07
DM BeLFusion 0.507 0.596 9.914 0.543 0.606 7.750 16.812 9.07 0.23 10.65 0.31
CoMusion 0.550 0.600 8.773 0.588 0.611 14.400 12.282 6.21 0.66 8.60 0.87
Ours 0.517 0.587 7.106 0.567 0.603 10.547 8.188 4.56 0.22 5.95 0.30
Table 10: Models trained on AMASS tested on zero-shot on 3DPW with synthetic noise up to 2cm added to 25% of the input.
Precision Body Realism
mean $\downarrow$ RMSE $\downarrow$
ADE $\downarrow$ FDE $\downarrow$ MAE $\downarrow$ str jit str jit
DLow 0.716 0.776 12.397 7.36 0.23 9.57 0.40
DivSamp 0.728 0.879 12.373 5.01 0.23 7.49 0.32
BeLFusion 0.657 0.756 11.175 8.89 0.18 10.69 0.27
CoMusion 0.670 0.792 10.215 4.56 0.33 6.28 0.46
SkeletonDiffusion 0.660 0.779 9.045 3.67 0.14 4.94 0.24
Table 11: Long term prediction (5s) on AMASS via autoregression of models traimed to predict 2s. MMGT is undefined in this case.
[…]
DLow 31 MB 8.1 M 111 ms
DivSamp 88 MB 23.1 M 8 ms
BeLFusion 53 MB 17.8 M 10 341 ms
HumanMAC 114 MB 28.7 M 7 438 ms
Comusion 87 MB 19 M 153 ms
SkeletonDiffusion 106 MB 26.5 M 412 ms
Table 12: Model footprint for a single H36M inference (RTX 6000)
F.1 Diversity and Body Realism
F.2 Human3.6M
F.3 Challenging Scenario: Synthetic Noise in Zero-Shot Generalization
We perform further experiments on the out-of-distribution, in-the-wild data of 3D Poses in the wild (3DPW), evaluated in Tab. 3, by designing a challenging scenario with synthetic noise (Tab. 10). We add random noise of a maximal magnitude of 2cm to 25% of the input observation keypoint, thus testing robustness to noise for models that were trained with precise, MoCap data (AMASS). While the experiments in Tab. 2 show models trained on noisy data (FreeMan), here we test robustness to noise in a zero-shot setting. SkeletonDiffusiondelivers among the highest precision and diversity, and the most realistic motions with a gap between 26% and 65% compared to the otherwise closest competitor, CoMusion (see Tab. 1). While BeLFusion shows jitter values close to ours, the limb stretching and the CMD are almost double as high, meaning that the length of their limbs highly varies over the whole prediction timespan, and the joint velocities are unrealistic: they achieve high precision with extremely unrealistic motions.
F.4 Long Term Prediction
We test models trained on AMASS to predict the next 2s in the generation of 5s motions via autoregression (Tab. 11). Here we focus on Precision and Realism, as the multimodal GT is ill-defined in this setting, and diversity evaluation loses meaning and its measuremnt is polluted by the difficulty of the task. We achieve again the highest realism and SoTA precision demonstrating the effectivness of our explicit bias on the human skeleton.
F.5 Computational Efficiency
Measurement are reported in Tab. 12. While there is no obvious computational difference between diffusion models in latent (BeLFusion, Ours) and input space (HumanMAC, CoMusion), latent models achieve much better body realism, particularly jitter (Tab. 1), by not working with 3D coordinates directly.